OpenStudy (anonymous):

Find the product. Matrices help! {{5, 5, -2, 3}}{{-4, 9, 8, 7}}

OpenStudy (anonymous):

is that a row matrix

OpenStudy (anonymous):

[-20,45,-16,21]

OpenStudy (anonymous):



OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[\left[\begin{matrix}5 & 5 \\ -2 & 3\end{matrix}\right]\left[\begin{matrix}-4 & 9 \\ 8 & 7\end{matrix}\right]=\left[\begin{matrix}(5*-4)+(5*8) & ? \\ ? & ?\end{matrix}\right]\]nd now you can do it

OpenStudy (anonymous):